# State Space Models are Strong Text Rerankers

### v1

@@ Abstract
This study benchmarks SSM-based architectures (specifically, Mamba-1 and Mamba-2) against transformer-based models across various scales, architectures, and pre-training objectives, focusing on performance and efficiency in text reranking tasks. We find that (1) Mamba architectures achieve competitive text ranking performance, comparable to transformer-based models of similar size; (2) they are less efficient in training and inference compared to transformers with flash attention; and (3) Mamba-2 outperforms Mamba-1 in both performance and efficiency. These results underscore the potential of state space models as a transformer alternative and highlight areas for improvement in future IR applications.
@@ 1 Introduction
The transformer architecture Vaswani et al. (2017) is an established standard within NLP and IR community. Compared to recurrent neural networks (RNNs) transformers better capture long-range dependencies and also admit large scale pre-training. However, for inference with a sequence of length $L$ and $D$-dimensional hidden states, transformers cost $O(L)$ time and $O(LD)$ space complexity — proving to be less efficient than RNNs
@@
In essence, state space models compress the context into a smaller state of size $N$, achieving $O(1)$ time complexity and $O(ND)$ space complexity in inference time. However, the capabilities of SSMs are limited by the amount of information that can be compressed in its hidden state. To mitigate this, Gu and Dao (2023) propose a novel selective state space model named Mamba. Mamba selectively encodes the input to the hidden state to improve model expressiveness, while also addressing the computation problem with a selective scan method and hardware-aware optimization. Gu and Dao (2023) and followup work (Dao and Gu, 2024; Zhu et al., 2024, inter alia) examine the efficacy of Mamba models for various sequence modeling tasks, notably language modeling, and also image and audio tasks. The parameterized SSMs are able to achieve performance close to transformer-based models of similar sizes while also demonstrating efficiency in training and inference.
@@
To this end, we conduct a rigorous benchmarking study comparing the two model families, across varying architectures, sizes, pre-training objectives, and attention patterns. Specifically, we train neural reranking models following established training methodologies outlined in prior literature Boytsov et al. (2022); Gao et al. (2021); Ma et al. (2023). Our experiments allow us to address the two research questions above. We find that:
@@
    Although Mamba architectures have better complexity theoretically, in practice they are less efficient compared to transformer architecture with IO-aware optimization (e.g., flash attention Dao (2024)).
@@ 3 The Text Reranking Problem
Modern IR systems employ a two-stage retrieval-reranking pipeline (Schütze et al., 2008; Zhang et al., 2021; Asai et al., 2024, inter alia). After the initial retrieval by an efficient, scalable first-stage retriever, a reranker refines the ranklist to optimize ranking metrics. Reranking involves ordering texts (passages or documents) by their relevance to a query, with passage reranking being a finer-grained form of document reranking. Our focus is to study this second stage and perform a comprehensive analysis of different rerankers for the tasks of both passage reranking and document reranking.
Let $q$ be an input query, and $d\in\mathcal{D}$ be a text, where $\mathcal{D}$ is the set of all texts (passages for passage reranking and documents for document reranking). The reranking model $f_{\theta}(q,d)$, parameterized by $\theta$, predicts a scalar relevance score. The model $f$ is instantiated as a linear layer on top of a language model. We adopt the common practice (Yates et al., 2021; Nogueira et al., 2019; Boytsov et al., 2022; Ma et al., 2023, inter alia) of concatenating the query and the document as input to the model.
@@ 4.1 Passage Reranking
Our code is implemented in PyTorch (Paszke et al., 2019) using the Huggingface library Wolf (2019). The weights of the pre-trained language models are obtained from the Huggingface Hub [^1] https://huggingface.co/models. Wherever applicable, we use Flash Attention 2 Dao (2024), gradient accumulation, and activation checkpointing. Note that the models are trained without parameter-efficient fine-tuning techniques such as LoRA Hu et al. (2021) which is different from Ma et al. (2023). We also do not investigate alternative compression techniques for improved parameter efficiency, such as low-rank factorization (Gupta et al., 2024), and leave these avenues for future research.
We do not extensively tune hyperparameters; as discussed by prior works Boytsov et al. (2022); Ma et al. (2023) fine-tuning of reranking models is less sensitive to hyperparameters. We found the vanilla AdamW optimizer along with learning rate warm-up with linear scheduler to work for all training runs. Our implementation as well as checkpoints are made public to facilitate reproducibility. [^2] https://github.com/zhichaoxu-shufe/RankMambaV2 Refer to Appx. D for an overview of hyperparameters throughout the experiments. Out implementation as well as checkpoints will be made public to facilitate reproducibility.
@@
4.2 Passage Reranking Results
In-domain Evaluation. We show the in-domain passage reranking results in Table 2. First, note that our trained models are comparable to previously reported results. For example, we report BERT-base achieves 38.5, 73.3, 73.1 on Dev, DL19, DL20 respectively, compared to MonoBERT Nogueira and Cho (2019)’s 37.2, 72.3 and 72.2. This suggests the correctness of our training setup.
Between transformer models of different architectures, we notice that encoder-only and encoder-decoder models outperform decoder-only models (OPT-125M and 350M in our case), despite OPT being pre-trained with more tokens. We hypothesize the reason is that the bi-directional attention in encoders better capture the interaction between query and document tokens. But decoder-only models are easier to scale.
Between transformer and Mamba architecture, Mamba models are able to achieve strong performance. For example, Mamba-2-370M achieves 38.6, 75.8 and 74.0 on three datasets compared to best transformer-based model BERT-large’s 39.1, 76.4 and 72.4, despite being uni-directional. The best transformer-based model—Llama-3.2-1B outperforms the Mamba models of similar sizes. However, note that Llama-3.2-1B is pre-trained on 15T tokens compared to Mamba model’s 300B tokens. We conclude that Mamba models are competitive in the passage ranking task.
Among Mamba models, we notice overall Mamba-2 achieves better performance than Mamba-1, while being pre-trained on the same number of tokens. A similar trend is shown in BEIR and document reranking results. This suggests Mamba-2 is a better SSM architecture compared to Mamba-1 for text reranking.
Out-of-domain Evaluation. We report part of BEIR results in Table 3 and leave full results to Appx. E. Overall, Mamba models are able to achieve competitive performance compared to transformer-based models of similar sizes. Specifically, Mamba-2-1.3B achieves 53.6 NDCG@10 averaged over 13 datasets compared to OPT-1.3B’s 52.7. Compared to baselines, Mamba-based models are only outperformed by the much larger RankLlama — a 7B model based on 7B-sized retrieval model RepLlama. This again verifies our findings in in-domain evaluation and suggests the efficacy of Mamba models in the passage reranking task.
One surprising observation is the underperformance of the Llama-3.2-1B model. This pre-trained model was not only trained on more tokens (15B) but was also trained on a much more diverse set of web documents.
@@ 4.3 Document Reranking
Among the baselines, we include MonoT5 Nogueira et al. (2020), RankT5 Zhuang et al. (2023) along with RankLlama that is the current state-of-the-art model using 7 Billion parameters. We also report two baseline runs from Boytsov et al. (2022): BERT-base-FirstP and BERT-base-MaxP as a sanity check for our implementation. MaxP method first segments the long document into several shorter passages, then uses the maximum relevance of segmented passages as the relevance of the document. We train document reranking models for each of the pre-trained models highlighted in § 4.1. Note that as encoder-only models have a fixed context length (e.g., BERT with 512), we do not have LongP variants for them.
@@ 4.5 Training Throughput and Inference Speed
As highlighted in prior research (Waleffe et al., 2024; Gu and Dao, 2023, inter alia), the true benefit of the Mamba models is realized with an improved inference speed. We do not observe this to be the case for the document reranking task (see Table 5). The main reason is for inference, reranking only requires one single forward computation compared to multiple forward computations in autoregressive generation.
@@ 5 Related Works
Fine-tuning pre-trained transformers has been the standard practice for text ranking tasks (Yates et al., 2021; Karpukhin et al., 2020, inter alia). Combining the query and the document as the input, the model predicts a scalar score indicating the relevance. Prior works have highlighted different aspects of training transformer-based text ranking models. Nogueira and Cho (2019); Nogueira et al. (2020); Dai and Callan (2019) are among the first efforts to showcase the effectiveness of fine-tuning pre-trained transformer-based language models. Gao et al. (2021) studied the retrieval-reranking pipeline and recommended training rerankers by sampling negatives from the results of first-stage retrievers. Li et al. (2023); Hofstätter et al. (2021) studied the effectiveness of chunking and pooling in long document ranking with shorter context transformer models. Boytsov et al. (2022) focused on benchmarking long context pre-trained transformers in long document ranking.
@@
Within the IR community, works have explored the possibility of using state space models as retriever Zhang et al. (2024) and reranker Xu (2024). This study extends prior works with more comprehensive experiments and points out new directions.
@@ 6 Conclusion and Future Work
We picture two future directions of this work: the task direction and the model direction. From the task perspective, the efficacy of state space models, including Mamba should be further examined in other IR tasks (e.g., text retrieval). From the model perspective, hybrid models (Lieber et al., 2024; Lenz et al., 2024; Glorioso et al., 2024, inter alia) have shown promise in certain NLP tasks. We believe the effectiveness of hybrid models should be thoroughly tested. Additionally, low-level optimization for state space models is an interesting challenge that may offer substantial improvements.
@@ Appendix D Hyperparameter Setting
We report the hyperparameters for passage reranking at Table 9 and document reranking at Table 10, respectively.
@@ Appendix E Full BEIR Results
Mamba-1-370M Mamba-2-370M Mamba-1-790M Mamba-2-780M OPT-1.3B LlaMA3.2-1B Mamba-1-1.4B Mamba-2-1.3B

### v3

@@ Abstract
This study benchmarks SSM-based architectures (specifically, Mamba-1 and Mamba-2) against transformer-based models across various scales, architectures, and pre-training objectives, focusing on performance and efficiency in text reranking tasks. We find that (1) Mamba architectures achieve competitive text ranking performance, comparable to transformer-based models of similar size; (2) they are less efficient in training and inference compared to transformers with flash attention; and (3) Mamba-2 outperforms Mamba-1 in both performance and efficiency. These results underscore the potential of state space models as a transformer alternative and highlight areas for improvement in future IR applications. [^1] The code for reproducing our experiments is available at https://github.com/zhichaoxu-shufe/RankMambaV2
@@ 1 Introduction
The transformer architecture Vaswani et al. (2017) is an established standard within NLP and IR community. Compared to recurrent neural networks (RNNs) transformers better capture long-range dependencies and also admit large scale pre-training. However, for inference with a sequence of length $L$ and $D$-dimensional hidden states, transformers cost $O(L)$ time and $O(LD)$ space complexity — proving to be less efficient than RNNs.
@@
In essence, state space models compress the context into a smaller state of size $N$, achieving $O(1)$ time complexity and $O(ND)$ space complexity in inference time. However, the capabilities of SSMs are limited by the amount of information that can be compressed in its hidden state. To mitigate this, Gu and Dao (2023) propose a novel selective state space model named Mamba. Mamba selectively encodes the input to the hidden state to improve model expressiveness, while also addressing the computation problem with a selective scan method and hardware-aware optimization. Gu and Dao (2023) and followup work (Dao and Gu, 2024; Zhu et al., 2024; Wang et al., 2024; Waleffe et al., 2024, inter alia) examine the efficacy of Mamba models for various sequence modeling tasks, notably language modeling, and also image and audio tasks. The parameterized SSMs are able to achieve performance close to transformer-based models of similar sizes while also demonstrating efficiency in training and inference.
@@
To this end, we conduct a rigorous benchmarking study comparing the two model families, across varying architectures, sizes, pre-training objectives, and attention patterns. Specifically, we train neural reranking models following established training methodologies outlined in prior literature Gao et al. (2021); Boytsov et al. (2022); Ma et al. (2023). Our experiments allow us to address the two research questions above. We find that:
@@
    Although Mamba architectures have better complexity theoretically, in practice they are less efficient compared to transformer architecture with I/O-aware optimization (e.g., flash attention Dao (2024)).
@@ 3 The Text Reranking Problem
Modern IR systems employ a two-stage retrieval-reranking pipeline (Schütze et al., 2008; Zhang et al., 2021; Asai et al., 2024; Xu et al., 2025, inter alia). After the initial retrieval by an efficient, scalable first-stage retriever, a reranker refines the ranklist to optimize ranking metrics. Reranking involves ordering texts (passages or documents) by their relevance to a query, with passage reranking being a finer-grained form of document reranking. Our focus is to study this second stage and perform a comprehensive analysis of different rerankers for the tasks of both passage reranking and document reranking.
Let $q$ be an input query, and $d\in\mathcal{D}$ be a text, where $\mathcal{D}$ is the set of all texts (passages for passage reranking and documents for document reranking). The reranking model $f_{\theta}(q,d)$, parameterized by $\theta$, predicts a scalar relevance score. The model $f$ is instantiated as a linear layer on top of a language model. We adopt the common practice of concatenating the query and the document as input to the model (Nogueira et al., 2019; Yates et al., 2021; Boytsov et al., 2022; Ma et al., 2023, inter alia).
@@ 4.1 Passage Reranking
Our code is implemented in PyTorch (Paszke et al., 2019) using the Huggingface library Wolf (2019). The weights of the pre-trained language models are obtained from the Huggingface Hub [^2] https://huggingface.co/models. Wherever applicable, we use Flash Attention 2 Dao (2024), gradient accumulation, and activation checkpointing. Note that the models are trained without parameter-efficient fine-tuning techniques such as LoRA Hu et al. (2021) which is different from Ma et al. (2023). We also do not investigate alternative compression techniques for improved parameter efficiency, such as low-rank factorization Gupta et al. (2024b), and leave these avenues for future research.
We do not extensively tune hyperparameters; as discussed by prior works Boytsov et al. (2022); Ma et al. (2023) fine-tuning of reranking models is less sensitive to hyperparameters. We found the vanilla AdamW optimizer along with learning rate warm-up with linear scheduler to work for all training runs. Refer to Appx. D for an overview of hyperparameters throughout the experiments. Out implementation as well as checkpoints will be made public to facilitate reproducibility.
@@
4.2 Passage Reranking Results
In-domain Evaluation. We show the in-domain passage reranking results in Table 2. First, note that our trained models are comparable to previously reported results. For example, we report that BERT-base achieves 38.5, 73.3, 73.1 on Dev, DL19, DL20 respectively, compared to MonoBERT Nogueira and Cho (2019)’s 37.2, 72.3 and 72.2. This suggests the correctness of our training setup.
Between transformer models of different architectures, we notice that both encoder-only and encoder-decoder models outperform decoder-only models (OPT-125M and 350M in our case), despite OPT being pre-trained with more tokens. We hypothesize the reason is that the bi-directional attention in encoders better capture the interaction between query and document tokens. But decoder-only models are easier to scale.
Between transformer and Mamba architectures, Mamba models are able to achieve strong performance. For example, despite being uni-directional, Mamba-2-370M achieves 38.6, 75.8, and 74.0 on three datasets compared to the best transformer-based model in that parameter range—BERT-large’s 39.1, 76.4, and 72.4. The overall best transformer-based model—Llama-3.2-1B outperforms the Mamba models of similar size. However, note that Llama-3.2-1B is pre-trained on 15T tokens compared to Mamba model’s 300B tokens. We conclude that Mamba models are competitive in the passage ranking task.
Among Mamba models, despite being trained on the same number of tokens, we notice overall that Mamba-2 achieves better performance than Mamba-1. A similar trend is shown in BEIR and document reranking results. In conclusion, Mamba-2 is a better SSM architecture compared to Mamba-1 for text reranking.
@@ 4.2 Passage Reranking Results
Out-of-domain Evaluation. We report part of BEIR results in Table 3 and leave full results to Appx. E. Overall, Mamba models are able to achieve competitive performance compared to the transformer-based models of similar sizes. Specifically, Mamba-2-1.3B achieves 53.6 NDCG@10 averaged over 13 datasets compared to OPT-1.3B’s 52.7. Compared to baselines, Mamba-based models are only outperformed by the much larger RankLlama — a 7B model based on 7B-sized retrieval model RepLlama. This reinforces our findings from in-domain evaluation, suggesting the efficacy of Mamba models in the passage reranking task.
One surprising observation is the underperformance of the Llama-3.2-1B model. This pre-trained model was not only trained on more tokens (15B) but was also trained on a much more diverse set of web documents. Ideally, a model pre-trained on a more diverse set of documents should perform better on out-of-domain evaluation sets, but we find that to not be the case with Llama-3.2-1B model.
@@ 4.3 Document Reranking
Among the baselines, we include MonoT5 Nogueira et al. (2020), RankT5 Zhuang et al. (2023) along with RankLlama that is the current state-of-the-art model using 7 Billion parameters. We also report two baseline runs from Boytsov et al. (2022): BERT-base-FirstP and BERT-base-MaxP as a sanity check for our implementation. MaxP method first segments the long document into several shorter passages, then uses the maximum relevance of segmented passages as the relevance of the document. We train document reranking models for each of the pre-trained models highlighted in § 4.1. Note that as encoder-only models have a fixed context length (ex: BERT with 512), we do not have LongP variants for them.
@@ 4.5 Training Throughput and Inference Speed
As highlighted in prior research (Waleffe et al., 2024; Gu and Dao, 2023, inter alia), the true benefit of the Mamba models is realized with an improved inference speed. We do not observe this to be the case for the document reranking task (see Table 5). The main reason is that for inference, reranking only requires one single forward computation compared to multiple forward computations in autoregressive generation. We further discuss the deficiency of Mamba models in § 4.6.
4.6 Profiling Inference Computation
(a) OPT-125M
(b) OPT-125M w/ FlashAttn
(c) Mamba-1-130M
(d) Mamba-2-130M
(e) OPT-350M
(f) OPT-350M w/ FlashAttn
(g) Mamba-1-370M
(h) Mamba-2-370M
Figure 2: Inference profiling results for Mamba models versus OPT models of similar size.
To better understand the inference performance of Mamba models, we use the PyTorch profiler [^4] https://pytorch.org/docs/stable/profiler.html to analyze the execution time of Mamba models at the operator level, comparing it to Transformer-based models of similar sizes. As in § 4.5, we use the DL19 document ranking evaluation set, an input length of 512, an evaluation batch size of 32, and the same hardware configuration. The results are presented in Fig. 2.
For Transformer-based models like OPT, I/O-related operators (e.g., aten::copy_, aten::to, aten::_to_copy, etc.) account for the majority of the execution time. Flash Attention Dao (2024) mitigates this by optimizing the I/O operations involved in attention computation, as seen in the reduced execution time for I/O-related operations in Figs. 2(b) and 2(f). This optimization leads to a noticeable speedup in inference, highlighting the importance of improving I/O efficiency for Transformers.
In contrast, the total execution time of Mamba-1 is dominated by operators such as aten::is_nonzero, aten::item, and aten::_local_scalar_dense. The first operator, aten::is_nonzero, checks whether tensors contain any non-zero elements, while aten::item and aten::_local_scalar_dense are used to extract scalar values from tensors. This suggests that Mamba-1’s architecture might suffer from computational inefficiencies due to an over-reliance on these scalar-extraction operations, which could be bottlenecking the performance. We hypothesize that these operations contribute to the model’s overall computational deficiency, particularly in comparison to models that utilize more efficient tensor operations. Mamba-2 improves upon this by parameterizing the Mamba-2 block, allowing for more effective utilization of matrix multiplication. This change is reflected in the elimination of the aforementioned scalar-extraction operators, with the new operator MambaSplitConv1D now accounting for over half of the total execution time. Mamba-2’s shift towards matrix multiplication suggests a more balanced computational load, although it still doesn’t fully close the gap in terms of inference speed compared to Transformer models with Flash Attention. This empirical evidence points to the need for further architectural refinement to optimize performance and better leverage compute-optimized hardware.
@@ 5 Related Works
Fine-tuning pre-trained transformers has been the standard practice for text ranking tasks (Yates et al., 2021; Karpukhin et al., 2020, inter alia). Combining the query and the document as the input, the model predicts a scalar score indicating the relevance. Prior works have highlighted different aspects of training transformer-based text ranking models. Nogueira and Cho (2019); Nogueira et al. (2020); Dai and Callan (2019) are among the first efforts to showcase the effectiveness of fine-tuning pre-trained transformer-based language models. Gao et al. (2021) studied the retrieval-reranking pipeline and recommended training rerankers by sampling negatives from the results of first-stage retrievers. Li et al. (2023); Hofstätter et al. (2021) studied the effectiveness of chunking and pooling in long document ranking with shorter context transformer models. Boytsov et al. (2022) focused on benchmarking long context pre-trained transformers in long document ranking. Refer to Yates et al. (2021); Xu et al. (2025) for detailed surveys.
@@
Within the IR community, works have explored the possibility of using state space models as retriever Zhang et al. (2024) and reranker Xu (2024). Our study extends prior works with more comprehensive experiments and points out new directions. Additionally, there is extensive literature on domain generalization and robustness analysis of different types of transformer models in NLP (Liang et al., 2023; Gupta et al., 2024a; Awadalla et al., 2022; Gupta et al., 2023). Our domain generalization analysis with different types of transformer and state-space models augments this line of work for text ranking.
@@ 6 Conclusion and Future Work
We picture two future directions of this work: the task direction and the model direction. From the task perspective, the efficacy of state space models, including Mamba should be further examined in other IR tasks (e.g., text retrieval). From the model perspective, hybrid models Lieber et al. (2024); Lenz et al. (2024); Glorioso et al. (2024); Nvidia (2025) have shown promise in certain NLP tasks. We believe the effectiveness of hybrid models should be thoroughly tested. Additionally, optimization for state space models is an interesting challenge that may offer substantial improvements.
@@ Appendix D Hyperparameter Setting
We show the hyperparameters in Table 9 and Table 10.
@@ Appendix E Full BEIR Results
Mamba-1-370M Mamba-2-370M Mamba-1-790M Mamba-2-780M OPT-1.3B Llama-3.2-1B Mamba-1-1.4B Mamba-2-1.3B